# LYT-NET: Lightweight YUV Transformer-based Network for Low-light Image Enhancement

### v5

@@ Abstract
This letter introduces LYT-Net, a novel lightweight transformer-based model for low-light image enhancement (LLIE). LYT-Net consists of several layers and detachable blocks, including our novel blocks—Channel-Wise Denoiser (CWD) and Multi-Stage Squeeze & Excite Fusion (MSEF)—along with the traditional Transformer block, Multi-Headed Self-Attention (MHSA). In our approach we adopt a dual-path approach, treating chrominance channels $U$ and $V$ and luminance channel $Y$ as separate entities to help the model better handle illumination adjustment and corruption restoration. Our comprehensive evaluation on established LLIE datasets demonstrates that, despite its low complexity, our model outperforms recent LLIE methods. The source code and pre-trained models are available at https://github.com/albrateanu/LYT-Net
@@
LLIE, transformer, deep learning
@@ I Introduction
Figure 1: Our model (LYT-Net) delivers state-of-the-art performance in LLIE task, while maintaining the best computational efficiency (results are plotted on LOL dataset [13]).
@@
In this letter, we propose a novel lightweight transformer-based approach called LYT-Net. Different from the existing transformer-based methods, our method focuses on computational efficiency while still producing state-of-the-art results. Specifically, we first separate chrominance from luminance employing the YUV color space. The chrominance information (channels $U$ and $V$) is initially processed through a specialized Channel-wise Denoiser (CWD) block, which reduces noise while preserving fine details. To minimize computational complexity, the luminance channel $Y$ undergoes convolution and pooling to extract features, which are subsequently enhanced by a traditional Multi-headed Self-Attention (MHSA) block. The enhanced channels are then recombined and processed through a novel Multi-stage Squeeze and Excite Fusion (MSEF) block. Finally, the chrominance channels $U$ and $V$ channels are concatenated with the luminance $Y$ channel and passed through a final set of convolutional layers to produce the restored image.
@@ III Results and Discussion
As shown in Table I, LYT-Net consistently outperforms the current state-of-the-art (SOTA) methods across all versions of the LOL dataset in terms of both PSNR and SSIM. Additionally, LYT-Net is highly efficient, requiring only 3.49G FLOPS and utilizing just 0.045M parameters, which gives it a significant advantage over other SOTA methods that are generally much more complex. The only exception is 3DLUT[35], which is comparable to our approach in terms of complexity. However, LYT-Net clearly surpasses the 3DLUT method in both PSNR and SSIM. This combination of strong performance and low complexity highlights the overall effectiveness of LYT-Net.
@@ V Conclusions
We introduce LYT-Net, an innovative lightweight transformer-based model for enhancing low-light images. Our approach utilizes a dual-path framework, processing chrominance and luminance separately to improve the model’s ability to manage illumination adjustments and restore corrupted regions. LYT-Net integrates multiple layers and modular blocks, including two unique components—Channel-Wise Denoiser (CWD) and Multi-Stage Squeeze & Excite Fusion (MSEF)—as well as the traditional Vision Transformer (ViT) block with Multi-Headed Self-Attention (MHSA). A comprehensive qualitative and quantitative analysis demonstrates that LYT-Net consistently outperforms state-of-the-art (SOTA) methods on all versions of the LOL dataset in terms of PSNR and SSIM, while maintaining high computational efficiency.

### v6

@@ Abstract
This letter introduces LYT-Net, a novel lightweight transformer-based model for low-light image enhancement. LYT-Net consists of several layers and detachable blocks, including our novel blocks—Channel-Wise Denoiser (CWD) and Multi-Stage Squeeze & Excite Fusion (MSEF)—along with the traditional Transformer block, Multi-Headed Self-Attention (MHSA). In our method we adopt a dual-path approach, treating chrominance channels $U$ and $V$ and luminance channel $Y$ as separate entities to help the model better handle illumination adjustment and corruption restoration. Our comprehensive evaluation on established LLIE datasets demonstrates that, despite its low complexity, our model outperforms recent LLIE methods. The source code and pre-trained models are available at https://github.com/albrateanu/LYT-Net
@@
Low-light Image Enhancement, Vision Transformer, Deep Learning
@@ I Introduction
Figure 1: Our model (LYT-Net) delivers SOTA performance in LLIE task, while maintaining the best computational efficiency (results are plotted on LOL dataset [13]).
@@
In this letter, we propose a novel lightweight transformer-based approach called LYT-Net. Different from the existing transformer-based methods, our method focuses on computational efficiency while still producing state-of-the-art (SOTA) results. Specifically, we first separate chrominance from luminance employing the YUV color space. The chrominance information (channels $U$ and $V$) is initially processed through a specialized Channel-wise Denoiser (CWD) block, which reduces noise while preserving fine details. To minimize computational complexity, the luminance channel $Y$ undergoes convolution and pooling to extract features, which are subsequently enhanced by a traditional Multi-headed Self-Attention (MHSA) block. The enhanced channels are then recombined and processed through a novel Multi-stage Squeeze and Excite Fusion (MSEF) block. Finally, the chrominance channels $U$ and $V$ channels are concatenated with the luminance $Y$ channel and passed through a final set of convolutional layers to produce the restored image.
@@ III Results and Discussion
As shown in Table I, LYT-Net consistently outperforms the current SOTA methods across all versions of the LOL dataset in terms of both PSNR and SSIM. Additionally, LYT-Net is highly efficient, requiring only 3.49G FLOPS and utilizing just 0.045M parameters, which gives it a significant advantage over other SOTA methods that are generally much more complex. The only exception is 3DLUT[35], which is comparable to our approach in terms of complexity. However, LYT-Net clearly surpasses the 3DLUT method in both PSNR and SSIM. This combination of strong performance and low complexity highlights the overall effectiveness of LYT-Net.
@@ V Conclusions
We introduce LYT-Net, an innovative lightweight transformer-based model for enhancing low-light images. Our approach utilizes a dual-path framework, processing chrominance and luminance separately to improve the model’s ability to manage illumination adjustments and restore corrupted regions. LYT-Net integrates multiple layers and modular blocks, including two unique components—Channel-Wise Denoiser (CWD) and Multi-Stage Squeeze & Excite Fusion (MSEF)—as well as the traditional Vision Transformer (ViT) block with Multi-Headed Self-Attention (MHSA). A comprehensive qualitative and quantitative analysis demonstrates that LYT-Net consistently outperforms SOTA methods on all versions of the LOL dataset in terms of PSNR and SSIM, while maintaining high computational efficiency.